OpenStudy (anonymous):

(4 + sqrt 50) - (3 - sqrt 8)

OpenStudy (anonymous):

@ kaleighdunn can you tell me if the solution is correct or not (4 + sqrt 50) - (3 - sqrt 8) =4 + sqrt 50 - (3 - sqrt 8) =4 + sqrt 50 - 3 + sqrt 8 =1 + sqrt 50 + sqrt 8 is this correct? ..yes or no?

OpenStudy (jiteshmeghwal9):

@mark_o. u did it very correct but u haven't completed the solution yet\[1+5\sqrt{2}+2\sqrt{2}\]\[1+7\sqrt{2}\]

OpenStudy (jiteshmeghwal9):

but u deserves a medal :)